OpenStudy (laronjaxon):

Which of the following is the solution to the equation 8(y – 3) = 2y y = 4.5 y = 1.5 y = –1.5 y = –4.5

OpenStudy (texaschic101):

8(y - 3) = 2y (distribute 8 through the parenthesis) 8y - 24 = 2y. (subtract 8y from both sides) 8y - 24 - 8y = 2y - 8y -24 = - 6y (divide by -6 on both sides to get y by itself) -24/-6 = -6y/-6 4 = y check... 8(y - 3) = 2y 8(4 - 3) = 2(4) 8(1) = 8 8 = 8 (correct) this doesn't seem to be an answer choice.....??? My answer is : y = 4